OpenStudy (anonymous):

Please help me . Solve 3x - 4 = 7x + 9 x ≈ -25.86 x ≈ -2.59 x ≈ -0.39 x ≈ -0.04

OpenStudy (anonymous):

x = - 13/4

jimthompson5910 (jim_thompson5910):

Chances are that the problem is really \[\Large 3^{x-4}=7^{x+9}\]

OpenStudy (anonymous):

yes its expontents sorry

OpenStudy (anonymous):

i know right me too Lol.

OpenStudy (anonymous):

this should help... http://www.wolframalpha.com/input/?i=3%5E%28x-4%29%3D7%5E%28x%2B9%29

OpenStudy (anonymous):

Ohhh, with exponents duuh.

jimthompson5910 (jim_thompson5910):

To solve this, use natural logarithms \[\Large 3^{x-4}=7^{x+9}\] \[\Large \ln\left(3^{x-4}\right)=\ln\left(7^{x+9}\right)\] \[\Large (x-4)\ln\left(3\right)=(x+9)\ln\left(7\right)\] \[\Large x\ln\left(3\right)-4\ln\left(3\right)=x\ln\left(7\right)+9\ln\left(7\right)\] \[\Large x\ln\left(3\right)-x\ln\left(7\right)=9\ln\left(7\right)+4\ln\left(3\right)\] \[\Large x\left(\ln\left(3\right)-\ln\left(7\right)\right)=9\ln\left(7\right)+4\ln\left(3\right)\] \[\Large x=\frac{9\ln\left(7\right)+4\ln\left(3\right)}{\ln\left(3\right)-\ln\left(7\right)}\] \[\Large x\approx-25.85589028\]

OpenStudy (anonymous):

Thank you so much . (:

jimthompson5910 (jim_thompson5910):

yw